OpenStudy (anonymous):

a dolphin jumped out of the water with an initial velocity of 32 feet per second. After how many seconds did the dolphin enter the water? Use h=-16t^2+vt+s to find the answer

OpenStudy (precal):

bump the problem to the top as soon as you are able to

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

4 seconds

OpenStudy (anonymous):

thanks could u help me with another similar one please

OpenStudy (precal):

can you state the equation that you used to find 4 seconds? thanks

OpenStudy (anonymous):

use simple formula v=u+at ,where v= final velocity ,u=initial velocity ,a=acceleration due to gravity 32ft/sec and t=time when dolphin jumps out then say at time t it reaches maximum height and its final velocity will be 0, so 0=u-32t u =32ft/sec so time to reach top =1second now same time to fall and reach water so total =2 seconds not 4

OpenStudy (precal):

so you did not use the equation given?

OpenStudy (anonymous):

ok thanks could you help me with a problem similar to that one please?

OpenStudy (anonymous):

i studied in SI unites as not from UK , so cant understand the equation , basically its some form of s=ut+ 1/2 ( at^2)

OpenStudy (precal):

@ewrtgsd are you not interested in the process at all? What will you do on an exam @uzmashahid15 Thanks, I know there are other methods.

OpenStudy (anonymous):

the equation which is given is just corollary by combining v^2=u^2 + 2as and s=ut+1/2 at ^2

OpenStudy (anonymous):

A airplane flies 4200 km east against the wind in 6 hours. The return flight withe wind takes 5 hours. The speed of the wind remains constant. Find the speed of the plane in still air and the speed of the wind

OpenStudy (anonymous):

i hv used directly v=u+at

OpenStudy (anonymous):

assume airplane speed as v and wind speed as w 1st case going against wind effective speed =v-w use distance =speed *time speed =distance /time so v-w=700 second case v+w =840 add both equations v=770kmph w=70kmph